Solve for x by factoring completely: x^4 - 1
Found 2 solutions by flame8855, lynnlo:
Answer by flame8855(424)

(x^2-1)(x^2+1)
x =+- 1
x^2+1=0
x^2=-1 = i^2
x = +-i

Answer by lynnlo(4176)

(x^4+1)(x^8-x^4+1)(x^2+1)(x^4-x^2+1)(x+1)(x^2-x+1)(x-1)(x^2+x+1)